OpenStudy (anonymous):

April adds the expressions 9x + 7 and 5x – 3. Evan adds the expressions 14x + 2 and 4x + 3. How much greater is Evan’s sum than April’s sum?

OpenStudy (one098):

Add them? o.o

OpenStudy (bohotness):

acctaly when you at them you subtract

OpenStudy (bohotness):

so you need to find what is x and much means subtract

OpenStudy (anonymous):

._.

OpenStudy (anonymous):

ok then

OpenStudy (bohotness):

[(14x + 2) + (4x + 3)] - [(9x + 7) + (5x - 3)] =

OpenStudy (bohotness):

what does that get you

OpenStudy (bohotness):

(18x + 5) - (14x + 4) = 18x + 5 - 14x - 4 = 4x + 1

OpenStudy (bohotness):

(18x + 5) - (14x + 4) = 18x + 5 - 14x - 4 = 4x + 1